# Solving Optimal Experiment Design with Mixed-Integer Convex Methods

### v6

@@ 3.2 Generalized self-concordance
By Carderera et al. (2021), the Frank-Wolfe algorithm also converges (with similar convergence rates to the $L$-smooth case) if the objective is generalized self-concordant.
Definition 3.7 (Generalized Self-Concordance Sun & Tran-Dinh (2019)).
@@
By Sun & Tran-Dinh (2019, Proposition 2), the composition of a generalized self-concordant function with a linear map is still generalized self-concordant. Hence, it suffices that we show that the functions $f(X)=-\log\det(X)$ and $g(X)=\operatorname{Tr}\left(X^{-p}\right)$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$, $p>0$, are generalized self-concordant for $X$ in the PD cone.
Notice that $f$ is the logarithmic barrier for the PSD cone which is known to be self-concordant (Nesterov & Nemirovskii, 1994). Convergence of Frank-Wolfe for the (D-Opt) Problem is therefore guaranteed. For convergence for the (GTI-Opt) Problem and (A-Opt) Problem, we can show that $g$ is self-concordant on a part of the PD cone.
Theorem 3.8.
@@
As with the convexity proofs in Section A.2, we bring our function back to the univariate case. We define $M(t)=V+tU$ where $V\in D$, $U$ a symmetric matrix and $t\in{\mathbb{R}}$ such that $V+tU\in\mathcal{D}$. By Lemma A.23, the second and third derivatives of $h(t)=\operatorname{Tr}\left(M(t)^{-p}\right)$ are given by
@@
11 11 $\leq\frac{M_{f}(p(p+1))^{1/2}}{(p+2)}\frac{1}{\sqrt[4]{\alpha^{2p}n}}$
@@
We conjecture that there is another value of $\nu$ with a tighter value for the constant $M_{f}$ satisfying the condition in Definition 3.7.
Corollary 3.9.
@@
The proof of Theorem 3.8 does not translate to the log variant of the A-criterion. Our computational experiments strongly suggest that the function $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ is generalized self-concordant.
@@ 3.3 Sharpness
The classical convergence speed of Frank-Wolfe methods given a convex objective is $\Theta(1/\epsilon)$ where $\epsilon$ is the additive primal error to the true optimum $x^{*}$, i.e. $f(\mathbf{x})-f(\mathbf{x}^{*})\leq\epsilon$. For the objectives of the problems defined in Section 2.1, we can prove a stronger property than convexity, namely sharpness on (a subset of) the feasible region $\mathcal{P}$. This improves the convergence rates of Frank-Wolfe, for the proof see Section C and Zhao (2025).
Definition 3.11 (Sharpness (Braun et al., 2022, Section 3.1.5)).
Let $\mathcal{X}$ be a compact convex set. A convex function $f$ is $(c,\theta)$-sharp (over $\mathcal{X}$) for $0<M<\infty$, $10<\theta<$ 1 if for all $x\in\mathcal{X}$ and $\mathbf{x}^{*}\in\Omega^{*}_{\mathcal{X}}$, we have
@@
To prove Theorem 3.12, we first prove that the functions $f(X)=-\log\det(X)$, $g(X)=\operatorname{Tr}(X^{-p})$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ with $p>0$, are strongly convex on a part of the PD cone.
Definition 3.13 (Strong Convexity).
@@
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma A.22, $f(X)$ is convex and the second derivative of $h$ is given by
@@
By using a similar idea to Lemma A.29, we find
@@
Let $A\in\mathcal{D}$ and $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. For the computation of $h^{\prime\prime}(t)$, see Lemma A.23.
@@
Since $A^{-p}$ is positive definite by Corollary A.19 and the trace is cyclic, we have
@@
Note that $BB$ is postive semi-definite. Thus, we can use the bound in Fang et al. (1994).
@@
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathcal{D}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma A.24, we have
@@
Note that the matrix $BB$ is PSD. The proof of Lemma A.29 can be adapted for one of matrices being symmetric and the other positive definite.
@@
By Lemma 3.10, we know that the largest eigenvalue of the information matrix can be bounded. As for the minimum eigenvalue, it can be lower bounded on
@@
so the convex hull of all domain-feasible integer points, see Figure 1. Any point $Y\in\mathcal{W}$ can be represented as a convex combination of the integer vertices $\{V_{i}\}\in\mathcal{W}$. By Lemma A.28, we have
@@
Proof of Theorem 3.12.
The functions of interest are compositions of strongly convex functions, see Lemma 3.14 and Lemma 3.15, and linear maps. Observe that the linear maps $X(\mathbf{x})$ and $X_{C}(\mathbf{x})$ are, in general, not injective and hence strong convexity is not preserved.
We can, however, follow the argumentation in Braun et al. (2022, Example 3.28). All of the functions of interest have the form $k(\mathbf{x})=r(B\mathbf{x})$ where $r$ is a strongly convex function. Notice that $B=A^{\intercal}*A^{\intercal}$ where * denotes the column-wise Kroenecker product. Observe that
@@
is the convex hull of the feasible integer points that are also in the domain. This is a polytope and we may use the bound in Braun et al. (2022, Lemma 3.27). Notice that the gradient of $k$ is given by $\nabla k(\mathbf{x})=B^{\intercal}\nabla r(B\mathbf{x})$. As in Definition 3.11 let $\Omega^{*}_{\mathcal{W}}$ denote the set of minimizers of $\min_{\mathbf{x}\in\mathcal{W}}k(\mathbf{x})$ and let $\mathbf{x}*\in\Omega^{*}_{\mathcal{W}}$. Then, we have $\left\langle\nabla k(\mathbf{x}*),\mathbf{x}-\mathbf{x}*\right\rangle\geq 0$ for $\mathbf{x}\in\mathcal{W}$.
@@
Since our region of interest $\mathcal{W}$ is a polytope Braun et al. (2022, Lemma 3.27) holds and there exists a constant $c>0$ such that the distance $\left\|B\mathbf{x}-B\mathbf{y}\right\|$ between any two points $\mathbf{x},\mathbf{y}\in\mathcal{W}$ is bounded by
@@
Let $C=A^{\intercal}$ and $B=\operatorname{diag}(\mathbf{x}-\mathbf{y})A$ and let us use the bound in Lemma A.30.
@@
We have thus established sharpness for all objective functions on $\mathcal{W}$. Observe that $\mathcal{W}=\mathcal{P}$ for the Fusion Problems, thus, linear convergence of Frank-Wolfe is guaranteed on them. In general, however, we have $\mathcal{P}\backslash\mathcal{W}\cap D\neq\emptyset$, see also Figure 1. That is there exists domain feasible points $\mathbf{y}\in\mathcal{P}$ which cannot be expressed as the convex combination of domain feasible integer points in $\mathcal{P}$. Hence, we have only linear convergence of Frank-Wolfe on the continuous subproblems if the corresponding continuous solution is in $\mathcal{W}$. In this case, we may use Theorem C.32 in Section C to argue linear convergence. For the Equation D-Opt Problem, we could adapt the proof in Zhao (2025) to the Blended Pairwise Conditional Gradient (BPCG) to argue linear convergence on all the subproblems, see Proposition C.36 in Section C. For the A-criterion and its log variant, we have only linear convergence on nodes where the continuous solution is in $\mathcal{W}$. We cannot guarantee that this is the case for all subproblems at the nodes.
(a) $A$-criterion
@@
From our experiments, we have empirical evidence that we have linear convergence of the dual gap for all Optimal Problems. In Figure 2, we show the trajectory of the dual gap at the root node for the $A$-criterion, $D$-criterion and $\log A$-criterion. After a burn-in phase, we see that the trajectory becomes parallel to $1/e^{ct}$ for some constant $c>0$ which is a representation of the linear convergence.
@@
Objective Convergence Linear Convergence
$f_{F}(\mathbf{x})=\log\det(X_{C}(\mathbf{x}))$checkmark $\checkmark$
$g_{F}(\mathbf{x})=\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})$checkmark $\checkmark$
$k_{F}(\mathbf{x})=\log\left(\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})\right)$checkmark $\checkmark$
$f(\mathbf{x})=-\log\det(X(\mathbf{x}))$checkmark $\checkmark$
$g(\mathbf{x})=\operatorname{Tr}(X(\mathbf{x})^{-p})$checkmark Not guaranteed
@@ 4 Solution methods
The new framework introduced in Hendrych et al. (2025b) is implemented in the Julia package Boscia.jl. It is a Branch-and-Bound (BnB) framework that utilizes Frank-Wolfe methods to solve the relaxations at the node level. The Frank-Wolfe algorithm (Frank & Wolfe, 1956; Braun et al., 2022), also called Conditional Gradient algorithm (Levitin & Polyak, 1966), and its variants are first-order methods solving problems of the type:
@@
where $f$ is a convex, Lipschitz-smooth function and $\mathcal{X}$ is a compact convex set. These methods are especially useful if linear minimization problems over $\mathcal{X}$ can be solved efficiently. The Frank-Wolfe methods used in Boscia.jl are implemented in the Julia package FrankWolfe.jl, see Besançon et al. (2022; 2025).
@@
One novel aspect of Boscia.jl is its use of a Bounded Mixed-Integer Linear Minimization Oracle (BLMO) as the Linear Minimization Oracle (LMO) in Frank-Wolfe. Typically, the BLMO is a MIP solver but it can also be a combinatorial solver. This leads to more expensive node evaluations but has the benefit that feasible integer points are found from the root node and the feasible region is much tighter than the continuous relaxation for many problems. In addition, Frank-Wolfe methods can be lazified, i.e. calling the LMO at each iteration in the node evaluation can be avoided, see (Braun et al., 2017).
@@
Observe that in contrast to the epigraph-based formulation approaches that generate many hyperplanes, this method works with the equivalent of a single supporting hyperplane given by the current gradient and moves this hyperplane until it achieves optimality. It is known that once the optimal solution is found, a single supporting hyperplane can be sufficient to prove optimality (described e.g. for generalized Benders in Sahinidis & Grossmann (1991)). Finding this final hyperplane, however, may require adding many hyperplanes beforehand at suboptimal iterations. In the case of the problems discussed in this paper, the constraint polytope is uni-modular. Adding hyperplanes created from the gradient will not maintain this structure and in consequence, yields a numerically more challenging MIP. Our approach, on the other hand, keeps the polytope and thereby its uni-modularity intact.
@@
For the line search problem within Frank-Wolfe, we utilize the Secant line search which has shown a superior performance for generalized self concordant functions, see Hendrych et al. (2025a).
@@
Outer Approximation schemes are a popular and well-established way of solving MINLPs (Kronqvist et al., 2019). This approach requires an epigraph formulation of (OEDP):
@@
This approach approximates the feasible region of (E-OEDP) with linear cuts derived from the gradient of the non-linear constraints, in our case $\nabla f$. Note that this requires the information matrix $X(\mathbf{x})$ at the current iterate $\mathbf{x}$ to be positive definite, otherwise an evaluation of the gradient is not possible or rather it will evaluate to $\infty$. The implementation is done with the Julia wrapper of SCIP, (Bestuzheva et al., 2021; 2023). Note that generating cuts that prohibit points leading to singular $X(\mathbf{x})$, we will refer to them as domain cuts, is non-trivial. Thus, we are applying this approach to the Fusion Problems only since their information matrix is always positive definite.
@@
Another Outer Approximation approach, as implemented in Pajarito.jl (Coey et al., 2020), represents the non-linearities as conic constraints. This is particularly convenient in combination with the conic interior point solver Hypatia.jl (Coey et al., 2022b) as it implements the $\log\det$ cone (the epigraph of the perspective function of $\log\det$) directly.
@@
For the representation of the trace inverse, we utilize the dual of the separable spectral function cone (Coey et al., 2022a, Section 6):
@@
For our purposes, $\mathcal{Q}$ is the PSD cone and the spectral function $\varphi$ is the negative square root whose convex conjugate is precisely the trace inverse, see (Coey et al., 2023, Table 1) and (Pilanci et al., 2015, Collary 1) [^4] For further details, see this discussion Kapelevich (2023)..
@@
Additionally, the conic formulation allows the computation of domain cuts for $\mathbf{x}$. Hence, this solver can be used on all problems. Note that we use HiGHS (Huangfu & Hall, 2018) as a MIP solver within Pajarito.jl.
@@
The conic solver from the previous approach was developed fairly recently. A previous conic approach was introduced first in Sagnol (2011) for the continuous case and later extended to the mixed-integer case in Sagnol & Harman (2015). They showed OEDP under the A-criterion and D-criterion can be formulated as a second-order cone program (SOCP). The SOCP model for (D-Opt) is given by
@@
where $Y$ is a $m\times n$ matrix and $\bm{\mu}$ is a $m\times 1$ vector. The problem size does not increase as much, with $m(2+n)$ variables and $23m+$ 2 constraints.
@@
The most general solver strategy for (OEDP) with matrix means criteria was introduced in Ahipaşaoğlu (2021). Like Boscia.jl, it is a Branch-and-Bound-based approach with a first-order method to solve the node problems. The first-order method in question is a coordinate-descent-like algorithm. As the termination criterion, this method exploits that the objective function is a matrix mean and shows the connection of the resulting optimization problem
@@
to the generalization of the Minimum Volume Enclosing Ellipsoid Problem (MVEP) (Ahipaşaoğlu, 2021). The variables $w$ can be interpreted as a probability distribution and the number of times the experiments are to be run is $wN$. Concerning (OEDP), one can say $x=wN$.
Note first that we have reimplemented the method in Julia and modified it so that we solve the formulation as depicted in (OEDP). Secondly, we have improved and adapted the step size rules within the first-order method, see Section B. Note further that the solver was developed for instances with a plethora of experiments and very few parameters. The solver employs the simplest Branch-and-Bound tree, i.e. with the most fractional branching rule and traverses the tree with respect to the minimum lower bound. In the next section, we will see that the method works well in cases where $n$ is small but struggles if the value of $n$ increases.

### v7

@@ 3.2 Generalized self-concordance
By carderera2021simple, the Frank-Wolfe algorithm also converges (with similar convergence rates to the $L$-smooth case) if the objective is generalized self-concordant.
Definition 4 (Generalized Self-Concordance sun2019generalized).
@@
By sun2019generalized, the composition of a generalized self-concordant function with a linear map is still generalized self-concordant. Hence, it suffices that we show that the functions $f(X)=-\log\det(X)$ and $g(X)=\operatorname{Tr}\left(X^{-p}\right)$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$, $p>0$, are generalized self-concordant for $X$ in the PD cone.
Notice that $f$ is the logarithmic barrier for the PSD cone which is known to be self-concordant (nesterov1994interior). Convergence of Frank-Wolfe for the (D-Opt) Problem is therefore guaranteed. For convergence for the (GTI-Opt) Problem and (A-Opt) Problem, we can show that $g$ is self-concordant on a part of the PD cone.
Theorem 4.
@@
As with the convexity proofs in Section˜A.2, we bring our function back to the univariate case. We define $M(t)=V+tU$ where $V\in D$, $U$ a symmetric matrix and $t\in{\mathbb{R}}$ such that $V+tU\in\mathcal{D}$. By Lemma˜10, the second and third derivatives of $h(t)=\operatorname{Tr}\left(M(t)^{-p}\right)$ are given by
@@
$1$ $\leq\frac{M_{f}(p(p+1))^{1/2}}{(p+2)}\frac{1}{\sqrt[4]{\alpha^{2p}n}}$
@@
We conjecture that there is another value of $\nu$ with a tighter value for the constant $M_{f}$ satisfying the condition in Definition˜4.
Corollary 1.
@@
The proof of Theorem˜4 does not translate to the log variant of the A-criterion. Our computational experiments strongly suggest that the function $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ is generalized self-concordant.
@@ 3.3 Sharpness
The classical convergence speed of Frank-Wolfe methods given a convex objective is $\Theta(1/\epsilon)$ where $\epsilon$ is the additive primal error to the true optimum $x^{*}$, i.e. $f(\mathbf{x})-f(\mathbf{x}^{*})\leq\epsilon$. For the objectives of the problems defined in Section˜2.1, we can prove a stronger property than convexity, namely sharpness on (a subset of) the feasible region $\mathcal{P}$. This improves the convergence rates of Frank-Wolfe, for the proof see Section˜C and zhao2025new.
Definition 5 (Sharpness (braun2022conditional, Section 3.1.5)).
Let $\mathcal{X}$ be a compact convex set. A convex function $f$ is $(c,\theta)$-sharp (over $\mathcal{X}$) for $0<M<\infty$, $0<\theta<1$ if for all $x\in\mathcal{X}$ and $\mathbf{x}^{*}\in\Omega^{*}_{\mathcal{X}}$, we have
@@
To prove Theorem˜5, we first prove that the functions $f(X)=-\log\det(X)$, $g(X)=\operatorname{Tr}(X^{-p})$ and $k(X)=\log\left(\operatorname{Tr}\left(X^{-p}\right)\right)$ with $p>0$, are strongly convex on a part of the PD cone.
Definition 6 (Strong Convexity).
@@
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Then we can define $h(t)=-\log\det\left(A+tB\right)$. By Lemma˜9, $f(X)$ is convex and the second derivative of $h$ is given by
@@
By using a similar idea to Lemma˜16, we find
@@
Let $A\in\mathcal{D}$ and $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ such that $A+tB\in\mathcal{D}$. Define $h(t)=\operatorname{Tr}((A+tB)^{-p})$. For the computation of $h^{\prime\prime}(t)$, see Lemma˜10.
@@
Since $A^{-p}$ is positive definite by Corollary˜2 and the trace is cyclic, we have
@@
Note that $BB$ is postive semi-definite. Thus, we can use the bound in fang1994inequalities.
@@
Let $A\in\mathcal{D}$, $B\in\mathbb{S}^{n}$ and $t\in{\mathbb{R}}$ s.t. $A+tB\in\mathcal{D}$. We define $h(t)=k(A+tB)=\log\left(\operatorname{Tr}((A+tB)^{-p})\right)$. By Lemma˜11, we have
@@
Note that the matrix $BB$ is PSD. The proof of Lemma˜16 can be adapted for one of matrices being symmetric and the other positive definite.
@@
By Lemma˜1, we know that the largest eigenvalue of the information matrix can be bounded. As for the minimum eigenvalue, it can be lower bounded on
@@
so the convex hull of all domain-feasible integer points, see Figure˜1. Any point $Y\in\mathcal{W}$ can be represented as a convex combination of the integer vertices $\{V_{i}\}\in\mathcal{W}$. By Lemma˜15, we have
@@
Proof of Theorem˜5.
The functions of interest are compositions of strongly convex functions, see Lemma˜2 and Lemma˜3, and linear maps. Observe that the linear maps $X(\mathbf{x})$ and $X_{C}(\mathbf{x})$ are, in general, not injective and hence strong convexity is not preserved.
We can, however, follow the argumentation in braun2022conditional. All of the functions of interest have the form $k(\mathbf{x})=r(B\mathbf{x})$ where $r$ is a strongly convex function. Notice that $B=A^{\intercal}*A^{\intercal}$ where * denotes the column-wise Kroenecker product. Observe that
@@
is the convex hull of the feasible integer points that are also in the domain. This is a polytope and we may use the bound in braun2022conditional. Notice that the gradient of $k$ is given by $\nabla k(\mathbf{x})=B^{\intercal}\nabla r(B\mathbf{x})$. As in Definition˜5 let $\Omega^{*}_{\mathcal{W}}$ denote the set of minimizers of $\min_{\mathbf{x}\in\mathcal{W}}k(\mathbf{x})$ and let $\mathbf{x}*\in\Omega^{*}_{\mathcal{W}}$. Then, we have $\left\langle\nabla k(\mathbf{x}*),\mathbf{x}-\mathbf{x}*\right\rangle\geq 0$ for $\mathbf{x}\in\mathcal{W}$.
@@
Since our region of interest $\mathcal{W}$ is a polytope braun2022conditional holds and there exists a constant $c>0$ such that the distance $\left\|B\mathbf{x}-B\mathbf{y}\right\|$ between any two points $\mathbf{x},\mathbf{y}\in\mathcal{W}$ is bounded by
@@
Let $C=A^{\intercal}$ and $B=\operatorname{diag}(\mathbf{x}-\mathbf{y})A$ and let us use the bound in Lemma˜17.
@@
We have thus established sharpness for all objective functions on $\mathcal{W}$. Observe that $\mathcal{W}=\mathcal{P}$ for the Fusion Problems, thus, linear convergence of Frank-Wolfe is guaranteed on them. In general, however, we have $\mathcal{P}\backslash\mathcal{W}\cap D\neq\emptyset$, see also Figure˜1. That is there exists domain feasible points $\mathbf{y}\in\mathcal{P}$ which cannot be expressed as the convex combination of domain feasible integer points in $\mathcal{P}$. Hence, we have only linear convergence of Frank-Wolfe on the continuous subproblems if the corresponding continuous solution is in $\mathcal{W}$. In this case, we may use Theorem˜6 in Section˜C to argue linear convergence. For the Equation˜D-Opt Problem, we could adapt the proof in zhao2025new to the Blended Pairwise Conditional Gradient (BPCG) to argue linear convergence on all the subproblems, see Proposition˜1 in Section˜C. For the A-criterion and its log variant, we have only linear convergence on nodes where the continuous solution is in $\mathcal{W}$. We cannot guarantee that this is the case for all subproblems at the nodes.
(a) AA-criterion
@@
From our experiments, we have empirical evidence that we have linear convergence of the dual gap for all Optimal Problems. In Figure˜2, we show the trajectory of the dual gap at the root node for the $A$-criterion, $D$-criterion and $\log A$-criterion. After a burn-in phase, we see that the trajectory becomes parallel to $1/e^{ct}$ for some constant $c>0$ which is a representation of the linear convergence.
@@
Objective
Convergence
Linear
Convergence
$f_{F}(\mathbf{x})=\log\det(X_{C}(\mathbf{x}))$ $\checkmark$ $\checkmark$
$g_{F}(\mathbf{x})=\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})$ $\checkmark$ $\checkmark$
$k_{F}(\mathbf{x})=\log\left(\operatorname{Tr}(X_{C}(\mathbf{x})^{-p})\right)$ $\checkmark$ $\checkmark$
$f(\mathbf{x})=-\log\det(X(\mathbf{x}))$ $\checkmark$ $\checkmark$
$g(\mathbf{x})=\operatorname{Tr}(X(\mathbf{x})^{-p})$ $\checkmark$ Not guaranteed
@@ 4 Solution methods
The new framework introduced in hendrych2023convex is implemented in the Julia package Boscia.jl. It is a Branch-and-Bound (BnB) framework that utilizes Frank-Wolfe methods to solve the relaxations at the node level. The Frank-Wolfe algorithm (frank1956algorithm; braun2022conditional), also called Conditional Gradient algorithm (levitin1966constrained), and its variants are first-order methods solving problems of the type:
@@
where $f$ is a convex, Lipschitz-smooth function and $\mathcal{X}$ is a compact convex set. These methods are especially useful if linear minimization problems over $\mathcal{X}$ can be solved efficiently. The Frank-Wolfe methods used in Boscia.jl are implemented in the Julia package FrankWolfe.jl, see besanccon2022frankwolfe; besanccon2025improved.
@@
One novel aspect of Boscia.jl is its use of a Bounded Mixed-Integer Linear Minimization Oracle (BLMO) as the Linear Minimization Oracle (LMO) in Frank-Wolfe. Typically, the BLMO is a MIP solver but it can also be a combinatorial solver. This leads to more expensive node evaluations but has the benefit that feasible integer points are found from the root node and the feasible region is much tighter than the continuous relaxation for many problems. In addition, Frank-Wolfe methods can be lazified, i.e. calling the LMO at each iteration in the node evaluation can be avoided, see (braun2017lazifying).
@@
Observe that in contrast to the epigraph-based formulation approaches that generate many hyperplanes, this method works with the equivalent of a single supporting hyperplane given by the current gradient and moves this hyperplane until it achieves optimality. It is known that once the optimal solution is found, a single supporting hyperplane can be sufficient to prove optimality (described e.g. for generalized Benders in sahinidis1991convergence). Finding this final hyperplane, however, may require adding many hyperplanes beforehand at suboptimal iterations. In the case of the problems discussed in this paper, the constraint polytope is uni-modular. Adding hyperplanes created from the gradient will not maintain this structure and in consequence, yields a numerically more challenging MIP. Our approach, on the other hand, keeps the polytope and thereby its uni-modularity intact.
@@
For the line search problem within Frank-Wolfe, we utilize the Secant line search which has shown a superior performance for generalized self concordant functions, see hendrych2025secant.
@@
Outer Approximation schemes are a popular and well-established way of solving MINLPs (kronqvist2019review). This approach requires an epigraph formulation of (OEDP):
@@
This approach approximates the feasible region of (E-OEDP) with linear cuts derived from the gradient of the non-linear constraints, in our case $\nabla f$. Note that this requires the information matrix $X(\mathbf{x})$ at the current iterate $\mathbf{x}$ to be positive definite, otherwise an evaluation of the gradient is not possible or rather it will evaluate to $\infty$. The implementation is done with the Julia wrapper of SCIP, (bestuzheva2021scip; bestuzheva2023enabling). Note that generating cuts that prohibit points leading to singular $X(\mathbf{x})$, we will refer to them as domain cuts, is non-trivial. Thus, we are applying this approach to the Fusion Problems only since their information matrix is always positive definite.
@@
Another Outer Approximation approach, as implemented in Pajarito.jl (CoeyLubinVielma2020), represents the non-linearities as conic constraints. This is particularly convenient in combination with the conic interior point solver Hypatia.jl (coey2022solving) as it implements the $\log\det$ cone (the epigraph of the perspective function of $\log\det$) directly.
@@
For the representation of the trace inverse, we utilize the dual of the separable spectral function cone (coey2022performance, Section 6):
@@
For our purposes, $\mathcal{Q}$ is the PSD cone and the spectral function $\varphi$ is the negative square root whose convex conjugate is precisely the trace inverse, see (coey2023conic, Table 1) and (pilanci2015sparse, Collary 1) [^4] For further details, see this discussion formulationforum..
@@
Additionally, the conic formulation allows the computation of domain cuts for $\mathbf{x}$. Hence, this solver can be used on all problems. Note that we use HiGHS (huangfu2018parallelizing) as a MIP solver within Pajarito.jl.
@@
The conic solver from the previous approach was developed fairly recently. A previous conic approach was introduced first in sagnol2011computing for the continuous case and later extended to the mixed-integer case in sagnol2015computing. They showed OEDP under the A-criterion and D-criterion can be formulated as a second-order cone program (SOCP). The SOCP model for (D-Opt) is given by
@@
where $Y$ is a $m\times n$ matrix and $\bm{\mu}$ is a $m\times 1$ vector. The problem size does not increase as much, with $m(2+n)$ variables and $3m+2$ constraints.
@@
The most general solver strategy for (OEDP) with matrix means criteria was introduced in ahipacsaouglu2021branch. Like Boscia.jl, it is a Branch-and-Bound-based approach with a first-order method to solve the node problems. The first-order method in question is a coordinate-descent-like algorithm. As the termination criterion, this method exploits that the objective function is a matrix mean and shows the connection of the resulting optimization problem
@@
to the generalization of the Minimum Volume Enclosing Ellipsoid Problem (MVEP) (ahipacsaouglu2021branch). The variables $w$ can be interpreted as a probability distribution and the number of times the experiments are to be run is $wN$. Concerning (OEDP), one can say $x=wN$.
Note first that we have reimplemented the method in Julia and modified it so that we solve the formulation as depicted in (OEDP). Secondly, we have improved and adapted the step size rules within the first-order method, see Section˜B. Note further that the solver was developed for instances with a plethora of experiments and very few parameters. The solver employs the simplest Branch-and-Bound tree, i.e. with the most fractional branching rule and traverses the tree with respect to the minimum lower bound. In the next section, we will see that the method works well in cases where $n$ is small but struggles if the value of $n$ increases.